# Generalized Face Liveness Detection via De-spoofing Face Generator

### v1

@@ 3.2 De-spoofing Face Generator
Figure 2: Structure of the Anomalous cue Guided FAS feature extraction Network (AG-Net). We take the residual of the input and the corresponding image generated by DFG as the anomalous cue, which guides the AG-Net to obtain a more robust FAS feature via the cross-attention module.
3.3 Anomalous Cue Guided Face Anti-spoofing
@@
After obtaining a well-trained DFG, we can generate a corresponding “real” face for every image in the FAS dataset. The residual between the generated “real” face and the input face is then taken as the anomalous cue which shows how different the input is from the real face distribution.
In some anomaly detection works [41, 44], the above residual can be directly used to calculate the anomaly scores for differentiating the normal and anomalous samples in the testing data. However, for the FAS task, this method leads to unsatisfactory performance, because the differences between real and spoof faces are relatively small and the result is easily affected by the background noise. Therefore, we resort to designing an Anomalous cue Guided FAS feature extraction Network (AG-Net), denoted as $\mathcal{F}_{AG}$, to learn FAS features with better generalization under a supervised training manner. As shown in Figure 2, the residual of the generated image and the input is first embedded by the anomalous cue encoder $\mathcal{F}_{cue}$. Meanwhile, the input is embedded by the ViT embedder $\mathcal{F}_{emb}$ [51] and put through the transformer-based model with $L$ cue guided feature extraction layers $\mathcal{F}_{guided}$. The process of the anomalous cue guided FAS feature extraction is as follows:
$f_{cue}=\mathcal{F}_{cue}(x-DFG(x)),$ (5)
$f_{0}=\mathcal{F}_{emb}(x),$ (6)
$f_{l}=\mathcal{F}_{guided}(f_{l-1},f_{cue}),\qquad l=1...L,$ (7)
$\mathcal{F}_{AG}(x)=f_{L}[0],$ (8)
where the first token of the final layer $f_{L}$ is taken as the extracted FAS feature of the AG-Net $\mathcal{F}_{AG}$.
At each cue guided feature extraction layer $\mathcal{F}_{guided}$, the anomalous cue embedding $f_{cue}$ is introduced to the mainstem feature extraction network and provides hints about the anomaly in the input via the cross-attention module inserted after the self-attention module [52]. The computation for the cue guided feature extraction layer is as follows:
$f_{l}^{\prime}=\mathrm{Attention}(f_{l}W_{Q}^{self},f_{l}W_{K}^{self},f_{l}W_{V}^{self}),$ (9)
$f_{l}^{\prime\prime}=\mathrm{Attention}(f_{l}^{\prime}W_{Q}^{cross},f_{cue}W_{K}^{cross},f_{cue}W_{V}^{cross}),$ (10)
$\mathcal{F}_{guided}(f_{l},f_{cue})=f_{l}^{\prime\prime}W_{foward},$ (11)
$\mathrm{Attention}(Q,K,V)=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,$ (12)
where $W_{Q}^{self},W_{K}^{self},W_{V}^{self}$ are learnable projection matrices in the self-attention module, $W_{Q}^{cross},W_{K}^{cross},W_{V}^{cross}$ are learnable projection matrices in the cross-attention module that introduces the anomalous cue guidance, and $W_{foward}$ is the parameter for the feed-forward calculation.
@@
As a Plug-and-Play feature extraction module, the proposed AG-Net can be combined with most previous DG-based FAS methods, further improving the generalization of the extracted FAS feature.
4 Experiments
4.1 Dataset
@@
We adopt four typical datasets to construct the most commonly used cross-domain benchmark: MSU-MFSD [2](denoted as M), CASIA-FASD [1](denoted as C), Idiap Replay-Attack [3](denoted as I), and OULU-NPU [4](denoted as O). Additionally, we introduce five other datasets that are commonly used in FAS evaluations: CelabA-Spoof [5], HKBU-MARs [6], WFFD [7], Rose-Youtu [9], SiW [8]. Based on all nine datasets, we conduct an experiment to further evaluate the generalization of our method with unseen scenarios and unknown presentation attacks. Some of the datasets above are in video form, and we uniformly sample 10 frames from each video in the time dimension for training and testing. For the training of the De-spoofing Face Generator (DFG), we use a private real face dataset that contains 100,000 unique face subjects and 200,000 real faces in total. Since this real face dataset involves privacy concerns, we cannot make it publicly available for now.
@@
4.2 Implementation Details
For the training of the DFG, we utilize the standard Latent Diffusion Model [48] as the backbone. The CLIP [50] textual encoder is replaced by the Arcface encoder [49] to get the face identity feature as the conditional guidance. The shape of the Arcface output before the pooling layer is $512\times 7\times 7$, which is reshaped to the size of $49\times 512$ and then padded as $49\times 768$ to fit the original conditional input of the U-Net in Latent Diffusion Model. During training. we set the batch size to 32 and train for a total of 150,000 steps on the real face dataset. The Adam optimizer is employed with a learning rate of 1e-05, a weight decay of 0.01, $\beta_{1}$ of 0.9, and $\beta_{2}$ of 0.999. The parameters of the Arcface model and the VQVAE in Latent Diffusion Model are kept fixed during training.
For the training of the FAS model, we take the ViT-base [51] model as the backbone of the Anomalous cue Guided FAS feature extraction Network (AG-Net). We then add the cross-attention modules used for introducing the anomalous cue after each self-attention module of the original ViT model. The input images are divided with a patch size of $16\times 16$ and embedded into image tokens in the shape of $196\times 768$. The anomalous cue of each image calculated by the DFG is first embedded by a cue encoder to the shape of $49\times 512$ and then put in the cross-attention modules of the AG-Net to provide hints about the anomaly. The cue encoder adopts the structure of a Resnet18 [54]. On the whole, we train our model using SSDG [19] strategy, and all hyper-parameters, like the learning rate and the optimizer, are set in accordance with the SSDG open-source code to ensure consistency and comparability.
Table 1: Evaluations on the Leave-One-Out Protocol
@@ 4.2 Implementation Details
AG-FAS (Ours) 5.71 97.76 5.89 98.14 5.86 98.36 9.41 96.36 6.72 97.65
[…]
4.3 Leave-One-Out Experiments

### v2

@@ III-A Overview
Figure 2: Structure of the Off-real Attention Network (OA-Net). We take the residual of the input and the corresponding image generated by DFG as the anomalous cue, which guides the OA-Net to obtain a more robust FAS feature via the cross-attention module.
III-B De-fake Face Generator
Firstly, we employ a dataset that contains only real faces to train a De-fake Face Generator (DFG), as shown in Figure 1, aiming to obtain a model that can generate corresponding “real” faces when given any faces as input. The real face dataset can be collected from the Internet or existing open-source facial datasets.
As for the backbone of our model, we choose Latent Diffusion Model (LDM) [58] for its excellent generation performance. We propose an identity branch to obtain the conditional input for the reverse process in order to enhance the consistency between the input and the generated face. Therefore, compared to using the LDM alone, our DFG can capture better anomalous cues, effectively reducing noise from inaccurate face reconstruction caused by identity mismatches between the reconstructed face and the input face. The training objective for DFG is as follows:
$L_{DFG}=\mathbb{E}_{x,E(x),\epsilon\in\mathcal{N}(0,1),t}\|\epsilon-\epsilon_{\theta}(z_{t},t,E_{id}(x))\|_{2}^{2},$ (4)
where $x$ is the input image, $E$ is the encoder which embeds the input $x$ into the clean latent code $z_{0}$, $z_{t}$ is the noisy latent code obtained via the forward process, $\epsilon$ is the Gaussian noise of the forward process, $\epsilon_{\theta}$ is the U-Net trained to predict the noise $\epsilon$, $t$ is the time step, and $E_{id}$ denotes the identity feature extractor. Specifically, an Arcface encoder [59] is used as the identity feature extractor to enhance the quality of the extracted features. During training, the parameters of the encoder $E$ and the identity feature extractor $E_{id}$ are fixed, and only the parameters of the U-Net $\epsilon_{\theta}$ are optimized.
Once the U-Net is trained, the DFG can be used to generate “real” faces corresponding to the given faces. Specifically, the DFG first embeds the input $x$ into the clean latent code $z_{0}$ and obtains the noisy $z_{\hat{t}}$ via a determined number of diffusion steps $\hat{t}$. Then, the U-Net takes the identity feature $E_{id}(x)$ of the input face as the conditional input and progressively generates the latent code $z_{0}^{\prime}$. Finally, we use the decoder $D$ in LDM to obtain the “real” face $x_{0}^{\prime}$ from the latent code $z_{0}^{\prime}$.
III-C Off-real Attention Network
@@
After obtaining a well-trained DFG, we generate a corresponding “real” face for every image in the FAS dataset. The residual between the generated “real” face and the input face is taken as the anomalous cue which shows how different the input is from the real face distribution.
We then design an Off-real Attention Network (OA-Net), as shown in Figure 2, to learn more robust FAS features with the help of the anomalous cue. The anomalous cue of the generated image is embedded by an anomalous cue encoder $E_{cue}$, utilizing a Resnet18 [60] backbone. Meanwhile, the original input is preprocessed and put into the ViT-based model with $L$ layers.
In order to exploit the anomalous cue, we follow the inspiration from the language model [61] and propose to add a cross-attention module in each layer of the model. When extracting the FAS feature, the OA-Net can obtain hints about the location of the off-real region by querying the embedded key and value of the anomalous cue at different layers. Compared to directly using ViT for feature extraction, focusing on off-real regions helps the OA-Net more effectively excavate robust FAS features. We initialize the weight of the last linear layer of the cross-attention module to be zero, so the mainstem network does not receive information from the anomalous cue branch at the beginning. The model will gradually learn to exploit the hints from the anomalous cue branch during the training process.
@@
Finally, the extracted FAS feature is fed into a classifier to distinguish between real and fake faces. The classifier is a linear model implemented with a fully connected layer. As a Plug-and-Play feature extraction module, the proposed OA-Net can be combined with most existing DG-based FAS methods, as shown in Figure 3, further improving the generalization of the extracted FAS feature.
Figure 3: The overall application diagram of AG-FAS, where OA-Net is integrated with existing DG-based FAS methods as a Plug-and-Play feature extraction module. $L_{DG}$ represents the feature constraint losses used in DG-based FAS methods, such as the adversarial loss and the asymmetric triplet loss in SSDG [21].
IV Theoretical Analysis
In this part, we provide theoretical guarantees for the distinguishability of the anomalous cues obtained from real and fake faces, which demonstrates the feasibility of the proposed framework. Considering the latent space learned by the autoencoding model in LDM [58] is perceptually equivalent to the image space, we also conduct our theoretical analysis in the latent space for simplification.
The theoretical analysis is performed in two directions: first, we prove the invariance of the latent codes for real faces before and after processing through the DFG, which indicates the corresponding anomalous cues tend to zero; then, we demonstrate that the output latent codes will differ from the fake input codes that contain spoof traces, resulting in noticeable anomalous cues for fake faces.
IV-A Anomalous Cue of Real Face
First of all, the following theorem reveals the invariance of the latent codes for real faces.
Theorem 1 For a latent code $z_{0}$ of a real face conforming to the distribution of the DFG training set and a given reconstruction error $\Delta z_{0}>0$, there exists a threshold time steps $t_{thr}$, such that when $t<t_{thr}$:
$\mathbb{E}_{z_{t}\in q(z_{t}|z_{0})}\|z_{0}-z_{0}^{\prime}\|_{2}^{2}<\Delta z_{0},$ (5)
where $q(z_{t}|z_{0})$ is the forward process, $z_{0}^{\prime}$ is the final output of the DFG reverse process using the sampled noisy latent code $z_{t}$.
Theorem 1 states that, by limiting the time steps of the forward process, the DFG model can effectively restore the original latent code $z_{0}$ based on the noisy latent code $z_{t}$ obtained from the forward process, which indicates the anomalous cue of a real input tends to be a matrix approaching $\mathbf{0}$.
Before introducing the proof for Theorem 1, we present the following assumption to ensure that the DFG model is sufficiently trained:
Assumption 1 We assume that the model is sufficiently trained. For a latent code $z_{0}$ of a real face conforming to the distribution of the DFG training set and a given predictive error $\Delta\epsilon>0$, there exists a threshold time step $t_{thr}$, such that when $t_{r}\leq t<t_{thr}$:
$\mathbb{E}_{z_{t}\in q(z_{t}|z_{0})}\|\epsilon-\epsilon_{\theta}(z_{t_{r}}^{\prime},t_{r},E_{id}(x))\|_{2}^{2}<\Delta\epsilon,$ (6)
where $q(z_{t}|z_{0})$ is the forward process, $z_{t}$ is the noisy latent code, $\epsilon$ is the Gaussian noise of the forward process, $z_{t_{r}}^{\prime}$ is the generated result in the reverse process at time step $t_{r}$, and $E_{id}(x)$ is the identity feature of the corresponding input.
The assumption implies that, within a certain time step, the well-trained model can effectively predict the added Gaussian noise $\epsilon$ in the forward process. With this assumption, we give the following proof of Theorem 1.
Proof of Theorem 1: Follow the definition in [62], the reverse process to generate $z_{t-1}^{\prime}$ from $z_{t}^{\prime}$ in the latent diffusion model is as follows:
$z_{t-1}^{\prime}$ $=\sqrt{\alpha_{t-1}}\left(\frac{z_{t}^{\prime}-\sqrt{1-\alpha_{t}}\cdot\epsilon_{\theta}^{t}}{\sqrt{\alpha_{t}}}\right)$ (7)
$+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{t}+\sigma_{t}\epsilon_{t},$
where $\epsilon_{\theta}^{t}$ represents $\epsilon_{\theta}(z_{t}^{\prime},t,E_{id}(x))$ for simplification, $\alpha_{0}$ is defined as 1, $\epsilon_{t}$ is a standard Gaussian noise, and $\sigma_{t}$ is a hyperparameter that is set as 0 in our DFG to obtain a denoising diffusion implicit model (DDIM) [62].
Thus, we can simplify the reverse process and express the generation output $z_{0}^{\prime}$ in a form represented by the noisy latent code $z_{t}$ as follows:
$z_{0}^{\prime}$ $=\sqrt{\alpha_{0}}\left(\frac{z_{1}^{\prime}-\sqrt{1-\alpha_{1}}\cdot\epsilon_{\theta}^{1}}{\sqrt{\alpha_{1}}}\right)+\sqrt{1-\alpha_{0}}\cdot\epsilon_{\theta}^{1}$ (8)
$=\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{1}}}z_{1}^{\prime}-\left(\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{1}}}\sqrt{1-\alpha_{1}}-\sqrt{1-\alpha_{0}}\right)\cdot\epsilon_{\theta}^{1}$
$=\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{2}}}z_{2}^{\prime}-\left(\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{2}}}\sqrt{1-\alpha_{2}}-\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{1}}}\sqrt{1-\alpha_{1}}\right)\cdot\epsilon_{\theta}^{2}$
$-\left(\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{1}}}\sqrt{1-\alpha_{1}}-\sqrt{1-\alpha_{0}}\right)\cdot\epsilon_{\theta}^{1}$
$=\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{t}}}z_{t}^{\prime}-\sum_{i=1}^{t}\left(\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{i}}}\sqrt{1-\alpha_{i}}\right.$
$\left.-\frac{\sqrt{\alpha_{0}}}{\sqrt{\alpha_{i-1}}}\sqrt{1-\alpha_{i-1}}\right)\cdot\epsilon_{\theta}^{i}.$
Since $\alpha_{0}$ is defined as 1 and $z_{t}^{\prime}$ is actually the noisy latent code $z_{t}$ obtained from the forward process, the generation output $z_{0}^{\prime}$ can be expressed as:
$z_{0}^{\prime}=\frac{1}{\sqrt{\alpha_{t}}}z_{t}-\sum_{i=1}^{t}\left(\frac{\sqrt{1-\alpha_{i}}}{\sqrt{\alpha_{i}}}-\frac{\sqrt{1-\alpha_{i-1}}}{\sqrt{\alpha_{i-1}}}\right)\cdot\epsilon_{\theta}^{i}.$ (9)
With the definition of the forward process $q(z_{t}|z_{0})$, the noisy latent code $z_{t}$ is obtained as:
$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$ (10)
where $\epsilon$ is the standard Gaussian noise used in the forward process. The expression of $z_{0}$ can then be represented by $z_{t}$ as follows:
$z_{0}=\frac{1}{\sqrt{\alpha_{t}}}z_{t}-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\epsilon.$ (11)
Taking the results from Eq. 9 and Eq. 11, we can calculate the reconstruction error for a given latent code $z_{0}$ of a real face as follows:
$\mathbb{E}_{z_{t}\in q(z_{t}|z_{0})}\|z_{0}-z_{0}^{\prime}\|_{2}^{2}$ (12)
$=\mathbb{E}_{z_{t}}\|\frac{1}{\sqrt{\alpha_{t}}}z_{t}-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\epsilon-\frac{1}{\sqrt{\alpha_{t}}}z_{t}$
$+\sum_{i=1}^{t}\left(\frac{\sqrt{1-\alpha_{i}}}{\sqrt{\alpha_{i}}}-\frac{\sqrt{1-\alpha_{i-1}}}{\sqrt{\alpha_{i-1}}}\right)\cdot\epsilon_{\theta}^{i}\|_{2}^{2}$
$=\mathbb{E}_{z_{t}}\|\sum_{i=1}^{t}\left(\frac{\sqrt{1-\alpha_{i}}}{\sqrt{\alpha_{i}}}-\frac{\sqrt{1-\alpha_{i-1}}}{\sqrt{\alpha_{i-1}}}\right)\cdot\left(\epsilon_{\theta}^{i}-\epsilon\right)\|_{2}^{2}$
$=\sum_{i=1}^{t}\left[\left(\frac{\sqrt{1-\alpha_{i}}}{\sqrt{\alpha_{i}}}-\frac{\sqrt{1-\alpha_{i-1}}}{\sqrt{\alpha_{i-1}}}\right)\mathbb{E}_{z_{t}}\|\epsilon_{\theta}^{i}-\epsilon\|_{2}^{2}\right].$
After applying Assumption 1, the upper bound of the reconstruction error turns out to be:
$\mathbb{E}_{z_{t}}\|z_{0}-z_{0}^{\prime}\|_{2}^{2}$ $<\sum_{i=1}^{t}\left(\frac{\sqrt{1-\alpha_{i}}}{\sqrt{\alpha_{i}}}-\frac{\sqrt{1-\alpha_{i-1}}}{\sqrt{\alpha_{i-1}}}\right)\Delta\epsilon$ (13)
$=\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\Delta\epsilon.$
Compared the upper bound with Eq. 5, we can conclude that Theorem 1 holds with a proper $t_{thr}$ that satisfies both the Assumption 1 and the following equation:
$\Delta\epsilon\leq\frac{\sqrt{\alpha_{t_{thr}}}}{\sqrt{1-\alpha_{t_{thr}}}}\Delta z_{0}<\frac{\sqrt{\alpha_{t}}}{\sqrt{1-\alpha_{t}}}\Delta z_{0}.$ (14)
It is worth noting that directly substituting the noise $\epsilon$ in Eq. 11 with the predicted result $\epsilon_{\theta}(z_{t},t,E_{id}(x))$ can also yield the upper bound in Eq. 13, which implies that a single-step reverse process can also obtain an upper bound on the reconstruction error.
IV-B Analysis for Forward Process
Before we turn to analyze the anomalous cues of fake faces, we need the following preliminary theorem.
Theorem 2 For the forward process of a latent diffusion model $q(z_{t}|z_{0})$, if the length $T$ of the forward process is sufficiently large, for a given $\epsilon_{KL}>0$ and two latent codes $z_{0}^{r},z_{0}^{f}$ of the real and fake faces from the same identity, there exists a time step $t$ such that:
$D_{KL}\{q(z_{t}^{r}|z_{0}^{r}),q(z_{t}^{f}|z_{0}^{f})\}<\epsilon_{KL},$ (15)
That is, the KL divergence between the latent distributions of the real and fake faces is sufficiently small under a certain time step $t$, making these two distributions indistinguishable.
The theorem indicates that the fake patterns in the latent codes can be removed through the forward process, which makes them indistinguishable from the latent codes of the real faces.
Proof of Theorem 2: Follow the definition in [62], the forward process of a latent diffusion model is as follows:
$q(z_{t}|z_{0}):=\mathcal{N}(z_{t};\sqrt{\alpha_{t}}z_{0},(1-\alpha_{t})I),$ (16)
where $\alpha_{1},...,\alpha_{T}$ represent a series of monotonically decreasing coefficients, ensuring that $10<\alpha_{T}<...<\alpha_{1}<$ 1.
The KL divergence between the latent distributions of the real and fake faces can then be computed:
$D_{KL}\{q(z_{t}^{r}|z_{0}^{r}),q(z_{t}^{f}|z_{0}^{f})\}$ (17)
$=D_{KL}\{\mathcal{N}(z_{t}^{r};\sqrt{\alpha_{t}}z_{0}^{r},(1-\alpha_{t})I),$
$\quad\quad\quad\quad\mathcal{N}(z_{t}^{f};\sqrt{\alpha_{t}}z_{0}^{f},(1-\alpha_{t})I)\}$
$=\log\frac{\sqrt{1-\alpha_{t}}}{\sqrt{1-\alpha_{t}}}+\frac{(1-\alpha_{t})+(\sqrt{\alpha_{t}}(z_{0}^{r}-z_{0}^{f}))^{2}}{2(1-\alpha_{t})}-\frac{1}{2}$
$=\frac{\alpha_{t}}{2(1-\alpha_{t})}(z_{0}^{r}-z_{0}^{f})^{2}.$
It can be observed that the KL divergence monotonically decreases as $\alpha_{t}$ decreases, and as $\alpha_{t}$ approaches 0, the KL divergence also approaches 0.
Taking the result into Eq. 15, we obtain the following condition required by Theorem 2:
$\frac{\alpha_{t}}{2(1-\alpha_{t})}(z_{0}^{r}-z_{0}^{f})^{2}<\epsilon_{KL}.$ (18)
After simplification, the range of the available value for $\alpha_{t}$ is as follows:
$\alpha_{t}<\frac{2\epsilon_{KL}}{2\epsilon_{KL}+(z_{0}^{r}-z_{0}^{f})^{2}}.$ (19)
Since the length $T$ is sufficiently large, $\alpha_{T}$ is set close to 0, which implies there exists a time step $t$ and its corresponding $\alpha_{t}$ such that:
$\alpha_{T}\leq\alpha_{t}<\frac{2\epsilon_{KL}}{2\epsilon_{KL}+(z_{0}^{r}-z_{0}^{f})^{2}},$ (20)
which means Theorem 2 holds.
The formula also shows that the smaller the difference between the two latent codes $z_{0}^{r}$ and $z_{0}^{f}$, the larger the upper limit for $\alpha_{t}$, and consequently, the smaller the time step $t$ is required in the forward process. When the two latent codes are equal, $\alpha_{1}<1$ with the time step $t=1$ just satisfies the condition.
IV-C Anomalous Cue of Fake Face
Theorem 2 shows that the forward process can remove the fake patterns, and Theorem 1 shows that the model can reconstruct the clean latent code of a real face from the noisy latent code. Combining Theorem 1 and Theorem 2, when the time step $t$ satisfies all the premises, we can obtain the following derivation for a fake input.
We first employ Theorem 2 and replace the noisy fake latent code $z_{t}^{f}$ with the noisy real latent code $z_{t}^{r}$:
$\mathbb{E}_{z_{t}^{f}\in q(z_{t}^{f}|z_{0}^{f})}\|z_{0}^{f}-z_{0}^{\prime}\|_{2}^{2}\approx\mathbb{E}_{z_{t}^{r}\in q(z_{t}^{r}|z_{0}^{r})}\|z_{0}^{f}-z_{0}^{\prime}\|_{2}^{2},$ (21)
where $z_{0}^{\prime}$ is the output of the reverse process determined by the sampled noisy latent code $z_{t}^{f}$ or $z_{t}^{r}$.
Then we utilize the triangle inequality and employ Theorem 1 to obtain the lower bound of the reconstruction error:
$\mathbb{E}_{z_{t}^{r}\in q(z_{t}^{r}|z_{0}^{r})}\|z_{0}^{f}-z_{0}^{\prime}\|_{2}^{2}$ (22)
$\geq\mathbb{E}_{z_{t}^{r}\in q(z_{t}^{r}|z_{0}^{r})}\left[\|z_{0}^{f}-z_{0}^{r}\|_{2}^{2}-\|z_{0}^{r}-z_{0}^{\prime}\|_{2}^{2}\right]$
$=\|z_{0}^{f}-z_{0}^{r}\|_{2}^{2}-\mathbb{E}_{z_{t}^{r}\in q(z_{t}^{r}|z_{0}^{r})}\|z_{0}^{r}-z_{0}^{\prime}\|_{2}^{2}$
$>\|z_{0}^{f}-z_{0}^{r}\|_{2}^{2}-\Delta z_{0},$
The derivation above indicates that when the input is a fake image, the residual between the input $z_{0}^{f}$ and the output $z_{0}^{\prime}$ of the DFG will reflect the input’s spoof traces $\|z_{0}^{f}-z_{0}^{r}\|_{2}^{2}$, resulting in a noticeable anomalous cue. Combining the result from Theorem 1, we prove the distinguishability between the anomalous cues obtained from real and fake faces.
IV-D Discussion
We find there exists a trade-off on the time step $t$ between Theorem 1 and Theorem 2: Theorem 1 requires a small time step to ensure the reconstruction of the latent code while Theorem 2 requires a large time step to remove the fake patterns.
The trade-off indicates that expanding the available range of the time step $t$ is the key to ensuring the practical performance of the DFG. Since the required $t$ in Theorem 2 is all determined by the difference between the input and its corresponding real face, we turn to study whether the threshold time step $t_{thr}$ in Theorem 1 can take a larger value. The proof of Theorem 1 indicates that, while keeping the reconstruction error of the latent code $\Delta z_{0}$ constant, increasing the value of the threshold time step $t_{thr}$ requires decreasing the prediction error of the noise $\Delta\epsilon$. Inspired by this observation, we utilize the identity feature as the guidance for the DFG model to obtain lower predictive error $\Delta\epsilon$, which allows us to choose a larger time step $\hat{t}$ for the DFG. Note that introducing the identity feature does not affect the derivation in Eq. 21 because the identity features are consistent within real and fake faces of the same people.
The ablation study of the trade-off on the time step $\hat{t}$ for the DFG is conducted in Sec. V-H3 and the results are shown in Table VI.
V Experiments
V-A Dataset
@@
We adopt four typical datasets to construct the most commonly used cross-domain benchmark: MSU-MFSD [3](denoted as M), CASIA-FASD [2](denoted as C), Idiap Replay-Attack [4](denoted as I), and OULU-NPU [5](denoted as O). Additionally, we introduce five other datasets that are commonly used in FAS evaluations: HKBU-MARs [7], WFFD [8], Rose-Youtu [10], SiW [9], WMCA [11]. Based on all nine datasets, we conduct an experiment to further evaluate the generalization of our method with unseen scenarios and unknown presentation attacks. Some of the datasets above are in video form, and we uniformly sample 10 frames from each video in the time dimension for training and testing.
For the training of the De-fake Face Generator (DFG), we use the data from FFHQ [63] and the real faces from CelebA-Spoof [6]. After filtering out images with deviated facial poses, we collect about 130,000 real faces in total.
@@
TABLE I: Evaluations on the Leave-One-Out Protocol
@@
DDM [54] 14.29 91.80 37.78 68.22 20.14 81.88 29.31 78.19 25.38 80.02
DDM* [54] 5.95 97.64 6.67 98.15 8.64 97.75 14.72 93.12 9.00 96.66
AG-FAS (Ours) 5.71 98.03 5.44 98.55 6.71 98.23 9.43 96.62 6.82 97.86
V-B Implementation Details
For the training of the DFG, we utilize the standard Latent Diffusion Model [58] as the backbone. Arcface encoder [59] is used to get the face identity feature as the conditional guidance. The shape of the Arcface output before the pooling layer is $512\times 7\times 7$, which is reshaped to the size of $49\times 512$ and then padded as $49\times 768$ to fit the original conditional input of the U-Net in Latent Diffusion Model. During training, we set the batch size to 32 and train for a total of 150,000 steps on the real face dataset. The Adam optimizer is employed with a learning rate of 1e-05, a weight decay of 0.01, $\beta_{1}$ of 0.9, and $\beta_{2}$ of 0.999. The maximum number of the diffusion steps is set to 1000. The parameters of the Arcface model and the encoder in Latent Diffusion Model are kept fixed during training.
For the training of the FAS model, we take the ViT-base [66] model as the backbone of the Off-real Attention Network (OA-Net). We then add the cross-attention modules used for introducing the anomalous cue after each self-attention module of the original ViT model. The input images are divided with a patch size of $16\times 16$ and embedded into image tokens in the shape of $196\times 768$. The anomalous cue of each image calculated by the DFG is first embedded by a cue encoder to the shape of $49\times 512$ and then put in the cross-attention modules of the OA-Net to provide hints about the anomaly. The cue encoder adopts the structure of a Resnet18 [60]. On the whole, we train our model using SSDG [21] strategy and use the open-source code to ensure consistency and comparability.
V-C Leave-One-Out Experiments